When a single card is drawn from a shuffled deck of cards, find the odds
against getting a 9.
1. 1:12
2. 12:1
3. 51:1
4. 1:51
I think it is 12:1
Please help!

Answer by stanbon(75887)

When a single card is drawn from a shuffled deck of cards, find the odds
against getting a 9.
1. 1:12
2. 12:1
3. 51:1
4. 1:51
I think it is 12:1
------------------
Odds against 9 = P(no 9)/P(9) = (12/13)/(1/13) = 12:1
====================
You are correct.
====================
Cheers,
Stan H.